OpenStudy (anonymous):

What is the first step in solving x2 + 2x + 3 = 0 by completing the square? A Factor the left side. B Add 3 to each side of the equation. C Divide both sides of the equation by 2 D Subtract 3 from each side of the equation.

OpenStudy (jhannybean):

\[x^2 +2x +3=0\] You want to group your like terms together. \[(x^2+2x)+3=0\] Now you will basically be completing the square inside the parenthesis. Tell me what you get when you have" (2/2)^2

OpenStudy (jhannybean):

Is that how you were told to solve i?

OpenStudy (jhannybean):

it*

OpenStudy (anonymous):

those are the answer choices, I think that it may either be A or B but im not sure how to solve this

OpenStudy (jhannybean):

Well, essentially when we "group" our like terms together, in a way we are completing the square by factoring, or creating a quadratic equation within the parenthesis.

OpenStudy (jhannybean):

B does notmake any sense.

OpenStudy (jhannybean):

The only time you add numbers to both sides of the equation is when you have fully factored what is inside the parenthesis.

OpenStudy (jhannybean):

Are you understanding?

OpenStudy (anonymous):

Yes now I see that adding 3 cannot be the solution and I think we have to factor it out before solving

OpenStudy (jhannybean):

B and C are random, unnecessary choices, I would go with either A or D.

OpenStudy (anonymous):

I agree I think that the answer is A

OpenStudy (jhannybean):

But I guess I am letting the way I complete the square interfere with the methods of completing the square are. so before factoring, you have to isolate your variables onto one side of the equation. how would you so that?

OpenStudy (jhannybean):

We want our x's on one side of the equation and our #s on the other.

OpenStudy (anonymous):

X^2+2X=-3

OpenStudy (jhannybean):

yes, and which answer choice do you think that is?

OpenStudy (anonymous):

Now I think that D is it because that should be done before factoring the equation

OpenStudy (jhannybean):

Yes! Good job :)

OpenStudy (jhannybean):

Did you want to also learn how to complete the square?

OpenStudy (anonymous):

Yes thank you so much! :)

OpenStudy (jhannybean):

Alright, so you have \[x^2 -2x = -3\] yes?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

in completing the square, we are trying to find a value of c that will put our equation in the form : \(ax^2 +bx+c=0\) We already have our \(ax^2+bx\). In order to find our \(c\) value, we shall take our coefficient of \(b\), divide it by 2, and square the entire quantity. We can see that in this format: \[c=\left(\frac{b}{2}\right)^2\]

OpenStudy (jhannybean):

So what is our value of b?

OpenStudy (anonymous):

2

OpenStudy (jhannybean):

try again :)

OpenStudy (jhannybean):

Remember, the equation you have can be written: \(x^2 +(-2x)=-3\)

OpenStudy (anonymous):

So it would be written as a negative?

OpenStudy (jhannybean):

yes :) -2 is our b value.

OpenStudy (jhannybean):

So what will we get when we plug it into that equation for finding c? \[c=\left(\frac{-2}{2}\right)^2\]

OpenStudy (jhannybean):

More specifically, what does c equal?

OpenStudy (anonymous):

1

OpenStudy (jhannybean):

Good :) c=1. Now we have completed the square on the left hand side, but one more thing you have to remember is, whatever you do to the left side of the equation, you must also do to the right side. we will have: \[x^2 -2x+1 = -3-1\]

OpenStudy (jhannybean):

Do you know what the left side of the equation can be reduced to?

OpenStudy (anonymous):

x^2-2x=-4

OpenStudy (jhannybean):

Not quite. I'll tell you a little trick.

OpenStudy (jhannybean):

When we were solving for c, we had reduced what was inside the parenthesis before squaring it, yes?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

So we had \[c=\left(\frac{-2}{2}\right)^2 = (-1)^2\] Taking that -1 value inside the square root, we can rewrite our equation on the left hand side as so: \[(x^2+c) =-4 \iff (x+(-1))^2 =-4 \iff (x-1)^2=-4\]

OpenStudy (jhannybean):

Let me know if you have any questions with this. It can be a little tricky but it's really useful

OpenStudy (anonymous):

I think I get this better now...That helped so much! :)

OpenStudy (jhannybean):

another way is to factor \((x^2-2x+1)\)

OpenStudy (jhannybean):

this might be a little easier to understand. Like to see?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Okay. What are 2 numbers that multiply to give +1 but add to give -2?

OpenStudy (anonymous):

-1

OpenStudy (jhannybean):

good :)

OpenStudy (jhannybean):

\[(x-1)(x-1)\]Since these are both the same terms, we can write them as \((x-1)^2\)

OpenStudy (jhannybean):

Does that make things more clear as to how i did it the first way?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Awesome :) Don't forget to medal the user who you think helped you understand :)

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (jhannybean):

no problem :) glad you understand!